# Chance-Constrained Covariance Steering for Discrete-Time Markov Jump Linear Systems

### v3

@@ Abstract
In this paper, we propose a novel convex optimization framework to solve the optimal covariance steering problem for discrete-time Markov Jump Linear Systems (MJLS) with chance constraints. We derive the analytical expressions for the mean and covariance trajectories of time-varying discrete-time MJLS and show that they cannot be separated even without chance constraints, unlike the single-mode dynamics case. To solve the covariance steering problem, we propose a two-step convex optimization framework, which optimizes the mean and covariance subproblems sequentially. Further, we use Gaussian approximations to incorporate chance constraints and propose an iterative optimization framework to solve the chance-constrained covariance steering problem. Both problems are originally nonconvex, and we derive convex relaxations which are proved to be lossless at optimality using the Karush–Kuhn–Tucker (KKT) conditions. Numerical simulations demonstrate the proposed method by achieving target covariances while respecting chance constraints under Gaussian noise and Markovian jump dynamics.
@@ I Introduction
Covariance steering is an important problem of designing a control policy that steers a system from an initial distribution to a target distribution while minimizing a cost function. It was first proposed in [1] as an infinite horizon problem and has attracted significant interest recently[2, 3, 4, 5].
@@
Despite recent remarkable advances in covariance steering theory for smooth dynamics, covariance steering for hybrid systems has been significantly underexplored. [13] studies covariance steering for a subclass of hybrid systems but don’t extend to Markov jump linear systems (MJLS). MJLS is a type of stochastic hybrid systems, where the mode dynamics are governed by a Markov chain. They have been extensively studied in the literature[14] with applications in fault-tolerant control [15], networks [16], and macroeconomic models [17].
Constrained control of MJLS has been an active area of research, and has benefited from the enormous progress in constrained control of single mode systems [bemporad_explicit_2002, 18] e.g. with state and control polytopic constraints for linear time-invariant[19] and periodic systems[20], and with second-moment constraints[21]. Chance constraints have previously been incorporated for MJLS[22], but they only consider time-invariant systems in a receding horizon framework using probabilistic reachable sets, which might not scale well with higher dimensions or longer time horizons. Meanwhile, the covariance steering paradigm offers a direct way to optimize for a given cost function under chance constraints in a computationally tractable manner. To the authors’ best knowledge, such formulations have previously not been investigated in the context of MJLS.
In this work, we focus on the optimal covariance steering problem for MJLS with chance constraints. This is motivated by the need for probabilistic path planning approaches which face abrupt changes in dynamics, such as spacecrafts[23] and robots[21] with sudden actuator failure. The contributions of this work are threefold: (1) Demonstrate the coupling of MJLS mean and covariance dynamics, showing why mean-covariance separation of single-mode systems does not extend to MJLS, (2) Derive a lossless convex relaxation to solve the MJLS covariance steering problem in a computationally efficient manner, and (3) Incorporate chance constraints in the covariance steering framework and propose a sequential convex programming approach to solve the problem.
@@
[…]
@@ II Problem Statement
We consider a discrete-time MJLS defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$ with a deterministic bias $c_{k}$ and a zero-mean Gaussian noise $w_{k}\sim\mathcal{N}(\mathbf{0},{I})$ given by:
@@
where $x\in\mathbb{R}^{n_{x}}$ and $u_{k}(i)\in\mathbb{R}^{n_{u}}$ are the state and control at time $k$ respectively. $r_{k}$ denotes the active mode of a homogeneous Markov chain taking values in $\Omega=\{1\ldots N\}$ with transition probability matrix $\mathbb{P}$ whose ($i$,$j$)th entry is $p_{ij}$. We define the following variables to study MJLS dynamics:
@@
$\mu_{k}\triangleq\mathbb{E}[x_{k}]=\sum_{i\in\Omega}q_{k}(i),$ $\Sigma_{k}\triangleq\mathbb{E}\left[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\right]$
The mode probability distribution vector is given by $\rho_{k}=[\rho_{k}(1)\ldots\rho_{k}(N)]$. For a given initial mode distribution $\rho_{0}$, $\rho_{k}$ can be calculated as $\rho_{k}=\rho_{0}\mathbb{P}^{k}$. We approximate $q_{k}(i)$ by $\rho_{k}(i)\mu_{k}$, which was found to be a reasonable approximation numerically. Similarly, we approximate $\mathbb{E}[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\mathds{1}_{r_{k}=i}]$ by $\rho_{k}(i)\Sigma_{k}$. The appendix section elaborates on this approximation and the consequent justification. We now define the MJLS covariance steering optimal control problem below:
@@
$x_{0}\sim\mathcal{N}(\overline{\mu}_{0},$ $\overline{\Sigma}_{0}),\;\mathbb{E}[x_{T}]=\mu_{f},\;\Sigma_{T}\preceq\Sigma_{f}$ (2c)
$P(x_{k}\in\mathcal{X})$ $\geq 1-\epsilon_{x},\quad P(u_{k}(i)\in\mathcal{U})\geq 1-\epsilon_{u}$ (2d)
where $Q_{k}\succeq 0$, $R_{k}\succ 0$ and $A_{k}(i)$ is invertible $\forall k\in\mathbb{Z}_{0:T},i\in\Omega$. The distribution $\rho_{0}$, $w_{k}$ and $x_{0}$ are independent.
@@
$\rho_{k}(i)>0,\;\forall k\in{\mathbb{Z}_{0:T}},i\in\Omega$, i.e. the probability of any mode being active is non-zero at all times.
@@
$\Sigma_{k}\succ 0$ $\forall k\in{\mathbb{Z}_{0:T}}$.
Problem 1 is hard to solve because of stochastic decision variables. To solve it in a computationally tractable manner, we focus on the mean and covariance propagation of the MJLS. We consider a mode-dependent affine feedback control policy:
Such control laws with a linear feedback law is common in covariance steering literature [11, 12] as it allows us to steer both the mean and covariance simultaneously. Focusing on the mean propagation first, we consider $q_{k}(i)$ below:
$q_{k+1}$ $(j)=\sum_{i\in\Omega}\mathbb{E}\left[\left({A}_{k}(i)x_{k}+B_{k}(i)(\overline{u}_{k}(i)+K_{k}(i)(x_{k}-\mu_{k}))\right.\right.$
@@
$=\sum_{i\in\Omega}p_{ij}\rho_{k}(i)\left({A}_{k}(i)\mu_{k}+B_{k}(i)\overline{u}_{k}(i)+c_{k}\right)$ (4)
Using the definition of $\mu_{k}$, eq. 4 can also be written as:
$\mu_{k+1}=\sum_{j\in\Omega}\sum_{i\in\Omega}p_{ij}\rho_{k}(i)\left({A}_{k}(i)\mu_{k}+B_{k}(i)\overline{u}_{k}(i)+c_{k}\right)$ (5)
Similarly, covariance propagation is studied below:
$\Sigma_{k+1}$ $=\left(\sum_{i\in\Omega}\mathbb{E}[(x_{k+1}x_{k+1}^{\top}\mathds{1}_{r_{k}=i}]\right)-\mu_{k+1}\mu_{k+1}^{\top}$ (6)
@@
${\Sigma}_{k+1}=\Bigl{(}\sum_{j\in\Omega}\sum_{i\in\Omega}\left[p_{ij}\rho_{k}(i)\{A_{k}(i){\Sigma}_{k}A^{\top}_{k}(i)\right.$
$\left.+A_{k}(i)(\mu_{k}\overline{u}^{\top}_{k}(i)+\Sigma_{k}K_{k}^{\top}(i))B^{\top}_{k}(i)+A_{k}(i)\mu_{k}c^{\top}_{k}\right.$
$\left.+B_{k}(i)(\mu_{k}\overline{u}^{\top}_{k}(i)+\Sigma_{k}K^{\top}_{k}(i))^{\top}A_{k}^{\top}(i)+B_{k}(i)\overline{u}^{\top}_{k}(i)c^{\top}_{k}\right.$
$\left.+B_{k}(i)(\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)+K_{k}(i)\Sigma_{k}K_{k}^{\top}(i))B_{k}^{\top}(i)+c_{k}c_{k}^{\top}\right.$
$\left.+c_{k}\mu^{\top}_{k}A_{k}^{\top}(i)+c_{k}\overline{u}^{\top}_{k}(i)B_{k}^{\top}(i)+G_{k}(i)G_{k}^{\top}(i)\right.$
$\left.+A_{k}(i)\mu_{k}\mu_{k}^{\top}A^{\top}_{k}(i)\}\right]\Bigr{)}-\mu_{k+1}\mu_{k+1}^{\top}$ (7)
@@
eq. 7 shows that the mean-covariance separation of single-mode dynamics ([10]) does not extend to MJLS.
We define a mean trajectory $\tau$ as a finite sequence of $(N+1)$-tuples $\{\mu_{k},\overline{u}_{k}(1)\ldots\overline{u}_{k}(N)\}_{k\in{\mathbb{Z}_{0:T-1}}}$ which satisfy eq. 5. We now define our optimal control problem in terms of $\mu_{k}$ and $\Sigma_{k}$, which approximates Problem 1.
@@
$\min_{\mu_{k},\overline{u}_{k}(i),K_{k}(i),\Sigma_{k}}\;J=\sum_{k=0}^{T-1}J_{k}$ (8)
$s.t.\;\forall k\in\mathbb{Z}_{0:T-1},$
@@
We write $J_{k}$ in terms of $\mu_{k}$ and $\Sigma_{k}$ as:
$J_{k}$ $=\sum_{i\in\Omega}\mathrm{Tr}\left(\mathbb{E}[x_{k}x_{k}^{\top}\mathds{1}_{r_{k}=i}]Q_{k}+\mathbb{E}[u_{k}(i)u^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]R_{k}\right)$
Using the identities derived in appendix A, we can rewrite $J_{k}$ as:
$J_{k}$ $=\sum_{i\in\Omega}\rho_{k}(i)\left[\mathrm{Tr}(\mu_{k}\mu_{k}^{\top}Q_{k})+\mathrm{Tr}(\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)R_{k})\right.$ (9)
$\left.\hskip 28.45274pt+\mathrm{Tr}({\Sigma}_{k}Q_{k})+\mathrm{Tr}(K_{k}(i)\Sigma_{k}K^{\top}_{k}(i)R_{k})\right]$
The nonlinearity in eq. 7 make Problem 2 results in a non-convex optimization landscape, while the chance constraints eq. 2d are not yet represented in closed form. We tackle these challenges in the following sections by first proposing a tight convex reformulation of Problem 2 without chance constraints, and then showing that these relaxations remain tight with a deterministic formulation of chance constraints too.

### v4

@@ Abstract
In this paper, we solve the chance-constrained covariance steering problem for discrete-time Markov Jump Linear Systems (MJLS) using a convex optimization framework. We derive the analytical expressions for the mean and covariance trajectories of time-varying discrete-time MJLS and show that they cannot be separated even without chance constraints, unlike the single-mode dynamics case. To solve the covariance steering problem, we propose a two-step convex optimization framework, which optimizes the mean and covariance subproblems sequentially. Further, we incorporate chance constraints and propose an iterative optimization framework to solve the chance-constrained covariance steering problem for MJLS. Both problems are originally nonconvex, and we derive convex relaxations which are proved to be lossless at optimality using the Karush–Kuhn–Tucker (KKT) conditions. Numerical simulations demonstrate the proposed method by achieving target covariances while respecting chance constraints under additive noise, bias, and Markovian jump dynamics.
@@ I Introduction
Covariance steering is an important problem of designing a control policy that steers a system from an initial given distribution to a target distribution while minimizing a cost function. It was first proposed in [1] as an infinite horizon covariance control problem, and has attracted significant interest recently as a finite horizon covariance steering problem[2, 3, 4, 5].
@@
Despite recent remarkable advances in covariance steering theory for smooth dynamics, covariance steering for hybrid systems has been significantly underexplored. [13] studies covariance steering for a class of hybrid systems that involve state-triggered switching. However, such approaches cannot extend to tackle covariance steering for Markov jump linear systems (MJLS). MJLS is a subclass of stochastic hybrid systems, where the mode dynamics are governed by a Markov chain. They have been extensively studied in the literature with applications in fault-tolerant control [14], network control [15], and macroeconomic models [16]. [17] is an excellent reference on the classical results for discrete-time MJLS.
Constrained control of MJLS has been an active area of research, e.g. with state and control polytopic constraints for linear time-invariant[18] and periodic systems[19], and with second-moment constraints[20]. Chance constraints have previously been incorporated for MJLS[21], but they only consider time-invariant systems in a receding horizon framework using probabilistic reachable sets, which might not scale well with higher dimensions or longer time horizons. Meanwhile, the covariance steering paradigm offers a direct way to optimize for a given cost function under chance constraints and boundary conditions in a computationally tractable manner. To the authors’ best knowledge, such formulations have previously not been investigated in the context of MJLS.
In this work, we focus on the optimal covariance steering problem for MJLS with chance constraints. This is motivated by the need for probabilistic motion planning approaches which face abrupt changes in dynamics, such as spacecrafts[22] and robots[20] with sudden actuator failure. The contributions of this work are threefold: (1) Demonstrate the coupling of MJLS mean and covariance dynamics, showing why mean-covariance separation of single-mode systems does not extend to MJLS, (2) Derive a lossless convex relaxation to solve the MJLS covariance steering problem in a computationally efficient manner, and (3) Incorporate chance constraints in the covariance steering framework and propose a sequential convex programming approach to solve the problem.
@@
$\mathbb{R}$ and $\mathbb{Z}_{a:b}$ denote the set of real numbers and integers from $a$ to $b$, respectively. $I$ denotes the identity matrix of appropriate size. $\mathds{1}_{\mathcal{\omega}}$ denotes the indicator function. $\lambda_{\max}(\cdot)$, $\mathrm{det}(\cdot)$, and $\mathrm{Tr(\cdot)}$ denote the maximum eigenvalue, determinant, and trace of a matrix respectively. $\|\cdot\|$ denotes the $l_{2}$ norm for a vector, while $\|\cdot\|_{p}$ denotes its $l_{p}$ norm.
@@ II Problem Statement
We consider a discrete-time MJLS defined on the probability space $(\Omega,\mathcal{F},\mathbb{P})$ with a wide-sense stationary white noise $w_{k}\in\mathbb{R}^{n_{w}}$ and a deterministic bias $c_{k}$, given by:
@@
where $x\in\mathbb{R}^{n_{x}}$ and $u_{k}(i)\in\mathbb{R}^{n_{u}}$ are the state and control at time $k$ respectively. $r_{k}$ denotes the active mode of a homogeneous Markov chain taking values in the set $\mathcal{S}\triangleq\{1\ldots N\}$ with transition probability matrix $\mathcal{P}$ whose (ii,$j$)th entry is $p_{ij}$. For the white noise $w_{k}$, we have $\mathbb{E}[w_{k}]=0$, $\mathbb{E}[w_{k}w^{\top}_{k}]=I_{n_{w}}$, and $\mathbb{E}[w_{j}w^{\top}_{k}]=0\;\forall i,j,k,j\neq k$ by definition. To study the MJLS dynamics, we focus on the mode probability $\rho_{k}(i)$, and the MJLS state mean $\mu_{k}$ and covariance $\Sigma_{k}$. In MJLS literature, it has typically been easier to work with partial expectations such as $\mathbb{E}[x_{k}\mathds{1}_{r_{k}=i}]$ instead of $\mathbb{E}[x_{k}]$ and we define such variables for the mean and covariance below:
@@
$\overline{x}_{k}(i)\triangleq\mathbb{E}[x_{k}|r_{k}=i],$ $\Sigma_{k}\triangleq\mathbb{E}\left[(x_{k}-\mu_{k})(x_{k}-\mu_{k})^{\top}\right]$
$\mu_{k}\triangleq\mathbb{E}[x_{k}],$ $\hskip-34.1433ptS_{k}(i)\triangleq\mathbb{E}[(x-\overline{x}_{k}(i))(x-\overline{x}_{k}(i))^{\top}\mathds{1}_{r_{k}=i}]$
The mode probability distribution vector is given by $\rho_{k}=[\rho_{k}(1)\ldots\rho_{k}(N)]$. For a given initial mode distribution $\rho_{0}$, $\rho_{k}$ can be calculated as $\rho_{k}=\rho_{0}\mathcal{P}^{k}$. Our framework can be extended trivially to time-inhomogeneous MJLS with $\rho_{k}=\rho_{0}\mathcal{P}_{0}\mathcal{P}_{1}\ldots\mathcal{P}_{k-1}$, but we only consider the time-homogeneous case for notational simplicity. The MJLS covariance steering optimal control problem is then defined as:
@@
$\mathbb{E}[x_{0}]=\mu_{0}$ $,\;\mathrm{cov}(x_{0})={\Sigma}_{0},\;\mathbb{E}[x_{T}]=\mu_{f},\;\mathrm{cov}(x_{T})\preceq\Sigma_{f}$ (2c)
$\mathbb{P}(x_{k}\in\mathcal{X})$ $\geq 1-\epsilon_{x},\quad\mathbb{P}(u_{k}(i)\in\mathcal{U})\geq 1-\epsilon_{u}$ (2d)
where $Q_{k}\succeq 0$, $R_{k}\succ 0$ and $A_{k}(i)$ is invertible $\forall k\in\mathbb{Z}_{0:T},i\in\mathcal{S}$. The distribution for $\rho_{0}$, $w_{k}$ and $x_{0}$ are independent. ${\mu}_{0},\;{\Sigma}_{0}$ are the initial mean and covariance, while $\mu_{f},\Sigma_{f}$ are the final time mean and covariance constraint with $\mathrm{cov}(x)$ denoting the covariance of $x$. $\mathcal{X}$ and $\mathcal{U}$ indicate the feasible sets which must be satisfied with probability $1-\epsilon_{x}$ and $1-\epsilon_{u}$ for state and control respectively.
@@
$\rho_{k}(i)>0,\;\forall k\in{\mathbb{Z}_{0:T}},i\in\mathcal{S}$, i.e. the probability of any mode being active is non-zero at all times.
Problem 1 is hard to solve as the decision variables are random variables. To solve it in a computationally tractable manner, we focus on the mean and covariance propagation of the MJLS. However, instead of dealing with $\mu_{k}$ and $\Sigma_{k}$, the MJLS dynamics motivate us to instead work with $q_{k}(i),\overline{x}_{k}(i)$ and $S_{k}(i)$, which are related to $\mu_{k}$ and $\Sigma_{k}$ as:
$\mu_{k}$ $=\sum_{i\in\mathcal{S}}q_{k}(i),\;\;\overline{x}_{k}(i)=\frac{q_{k}(i)}{\rho_{k}(i)}$ (3a)
$\Sigma_{k}$ $=\sum_{i\in\mathcal{S}}\left(S_{k}(i)+\rho_{k}(i)(\overline{x}_{k}(i)-\mu_{k})(\overline{x}_{k}(i)-\mu_{k})^{\top}\right)$ (3b)
We consider a mode-dependent affine feedback control policy:
@@
Under eq. 4, $S_{k}(i)\succ 0,\;\forall k\in{\mathbb{Z}_{0:T},i\in\mathcal{S}}$.
Such control laws with a linear feedback law is common in covariance steering literature [11, 12] as it allows us to steer both the mean and covariance simultaneously. We will now use the first and second order moments to solve Problem 1 in a tractable manner. Focusing on the mean, $q_{k}(i)$ propagation can be performed as given below:
$q_{k+1}$ $(j)=\mathbb{E}\left[x_{k+1}\mathds{1}_{\{r_{k+1}=j\}}\right]$
[…]
@@
$\hskip-14.22636pt={\sum_{i\in\mathcal{S}}p_{ij}\left({A}_{k}(i)q_{k}(i)+\rho_{k}(i)(B_{k}(i)\overline{u}_{k}(i)+c_{k})\right)}$ (5)
Similarly, covariance propagation can be studied using $S_{k}(i)$ as follows:
$S_{k+1}(j)$ $=\mathbb{E}[(x_{k+1}-\overline{x}_{k+1}(j))(x_{k+1}-\overline{x}_{k+1}(j))^{\top}\mathds{1}_{r_{k+1}=j}]$
$\hskip-14.22636pt=\mathbb{E}[x_{k+1}x_{k+1}^{\top}\mathds{1}_{r_{k+1}=j}]-\rho_{k+1}(j)\overline{x}_{k+1}(j)\overline{x}_{k+1}^{\top}(j)$ (6)
@@
$S_{k+1}(j)=\Bigl(\sum_{i\in\mathcal{S}}\left[p_{ij}\{A_{k}(i)S_{k}(i)A^{\top}_{k}(i)\right$.
$+A_{k}(i)S_{k}(i)K_{k}^{\top}(i)B_{k}^{\top}(i)+B_{k}(i)K_{k}(i)S_{k}^{\top}(i)A_{k}^{\top}(i)$
$+B_{k}(i)(\rho_{k}(i)\overline{u}_{k}(i)\overline{u}_{k}^{\top}(i)+K_{k}(i)S_{k}(i)K_{k}^{\top}(i))B_{k}^{\top}(i)$
$+\rho_{k}(i)(B_{k}(i)\overline{u}_{k}(i)c_{k}^{\top}+B_{k}(i)\overline{u}_{k}(i)\overline{x}_{k}^{\top}A_{k}^{\top}(i)+c_{k}c_{k}^{\top}$
$+c_{k}\overline{u}_{k}^{\top}(i)B_{k}^{\top}(i)+c_{k}\overline{x}_{k}^{\top}(i)A_{k}^{\top}(i)+A_{k}(i)\overline{x}_{k}(i)\overline{u}_{k}^{\top}(i)B_{k}^{\top}(i)$
$+A_{k}(i)\overline{x}_{k}(i)c_{k}^{\top}+A_{k}(i)\overline{x}_{k}(i)\overline{x}_{k}^{\top}(i)A_{k}^{\top}(i)+G_{k}(i)G_{k}^{\top}(i))\}\Bigr)$
$\hskip 71.13188pt-\rho_{k+1}(j)\overline{x}_{k+1}(j)\overline{x}_{k+1}^{\top}(j)$ (7)
@@
A careful analysis of eq. 7 along with eq. 3b show that the terms involving the MJLS mean information do not cancel out for the covariance propagation, demonstrating that the mean-covariance separation of single-mode dynamics (shown in [10]) does not extend to MJLS.
For a given finite sequence of tuples $\{q_{k}(1)\ldots q_{k}(N)\}_{k\in\mathbb{Z}_{0:T}}$ and $\{\overline{u}_{k}(1)\ldots\overline{u}_{k}(N)\}_{k\in\mathbb{Z}_{0:T-1}}$ which satisfy eq. 5, we define a corresponding mean trajectory $\tau$ as a set of two finite sequences of $N+1$ and $N$-tuples $\{\mu_{k},\overline{x}_{k}(1)\ldots\overline{x}_{k}(N)\}_{k\in\mathbb{Z}_{0:{T}}}$ $\{\overline{u}_{k}(1)\ldots\overline{u}_{k}(N)\}_{k\in\mathbb{Z}_{0:{T-1}}}$ respectively which satisfy eq. 3a and eq. 5 simultaneously.
We now redefine the cost function $J$ in terms of the newly defined mean and covariance variables. From eq. 2a, we know $J=\sum_{k=0}^{T-1}J_{k}$, where $J_{k}$ is given by
$J_{k}$ $=\sum_{i\in\mathcal{S}}\mathbb{E}\left[(x_{k}^{\top}Q_{k}x_{k}+u^{\top}_{k}(i)R_{k}u_{k}(i))\mathds{1}_{r_{k}=i}\right]$
$=\sum_{i\in\mathcal{S}}\mathrm{Tr}\left(\mathbb{E}[x_{k}x_{k}^{\top}\mathds{1}_{r_{k}=i}]Q_{k}+\mathbb{E}[u_{k}(i)u^{\top}_{k}(i)\mathds{1}_{r_{k}=i}]R_{k}\right)$
Using the identities derived in appendices A and A in the appendix, rewrite $J_{k}$ as:
$J_{k}$ $=\sum_{i\in\mathcal{S}}\left[\rho_{k}(i)\{\mathrm{Tr}(\overline{x}_{k}(i)\overline{x}_{k}^{\top}(i)Q_{k})+\mathrm{Tr}(\overline{u}_{k}(i)\overline{u}^{\top}_{k}(i)R_{k})\}\right$.
$\left.\hskip 14.22636pt+\mathrm{Tr}({S}_{k}(i)Q_{k})+\mathrm{Tr}(K_{k}(i)S_{k}(i)K^{\top}_{k}(i)R_{k})\right]$ (8)
We now define our optimal control problem in terms of the mean and covariance variables using eq. 4, which approximates Problem 1.
@@
$\min_{\mu_{k},q_{k}(i),\overline{x}_{k}(i),\overline{u}_{k}(i),K_{k}(i),S_{k}(i),\Sigma_{k}}\;J=\sum_{k=0}^{T-1}J_{k}$ (9)
$\mathrm{s.t.}\;\forall k\in\mathbb{Z}_{0:T-1},$
@@
Problem 2 is still not a tractable problem to solve because of the nonlinear constraint eq. 7 and chance constraints eq. 2d, which are not yet defined as a deterministic constraint in the mean and covariance variables. We tackle these challenges in the following sections by first proposing a lossless convex reformulation of Problem 2 without chance constraints, and then showing that these relaxations remain lossless with the deterministic formulations of chance constraints. Our lossless relaxations convert an equality constraint to an inequality constraint, which is then shown to be tight, i.e. equivalent to the original equality constraint at optimality.